OpenStudy (anonymous):

Can someone please help me out with this question!!! I have no idea where to start. The question is stated below: A typical industrial windmill has rotating blades that are each 116 ft long. From a distance, you observe that it takes a single blade 3.5 seconds to make a full rotation. How fast is the outer tip of the blade moving? Give your final answer in miles per hour.

jimthompson5910 (jim_thompson5910):

The tip makes a circle as it rotates. What is the circumference of this circle?

OpenStudy (anonymous):

728.48

jimthompson5910 (jim_thompson5910):

yes roughly, so it travels approximately 728.48 ft in 3.5 seconds

jimthompson5910 (jim_thompson5910):

3.5 seconds = 3.5/3600 = 0.000972 hours 728.48 ft = 728.48*(1/5280) = 0.1379697 miles "the tip travels approximately 728.48 ft in 3.5 seconds" is equivalent to saying "the tip travels approximately 0.1379697 miles in 0.000972 hours"

jimthompson5910 (jim_thompson5910):

Divide the two (distance/time) to get the speed in mph make sure you use the miles and hours

OpenStudy (anonymous):

why do you divide 3.5/3600

jimthompson5910 (jim_thompson5910):

because 1 hr = 60 min 1 min = 60 seconds so 1 hr = 60*60 = 3600 seconds

jimthompson5910 (jim_thompson5910):

3.5 seconds turns into 0.000972 hours (roughly) when you use the conversion 1 hr = 3600 seconds

OpenStudy (anonymous):

so would my answer be 141.9441358 miles per hour

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

yw